A stock currently sells for $50. In six months it will either rise to $60 or decline to $45. The continuous compounding risk-free interest rate is 5% per year.

  1. Using the binomial approach, find the value of a European call option with an exercise price of $50.
  2. Using the binomial approach, find the value of a European put option with an exercise price of $50.
  3. Verify the put-call parity using the results of Questions 1 and 2.
Answer #1

s=so, E= so, us = to, ds = 45 puo d = 45 S = 50 9 R = 50 -45 CEO 16=5 St., T = 6/6 ie 5 B us x 5 .025 103.53 Step-2 Risk neut